Amarindo, Inc is a newly public firm with 8.0 million shares outstanding. You are doing a valuation analysis of AMR. You estimate its free cash flow in the coming year to be 15.4 million, and you expect the firms free cash flows only to grow by 4.2% per year in subsequent years. Because the firm has only been listed on the stock exchange for a short time, you do not have an accurate assessment of AMR's equity beta. However, you do have beta data for UAL, another firm in the same industry. AMR has a much lower debt-equity ratio of 0.42, which is expected to remain stable, and its debt is risk-free, AMR's corporate tax is 30%, the risk-free rate is 5.2% and the expected return on the market portfolio is 11.2%

Debt Equity Ratio \= 1.4 Equity Beta= 2.10 Debt Beta=0.42

a. Estimate AMR's equity cost of capital

b, Estimate AMR's share price
Answer #1

a. First we calculate the weighted beta

First we calculate the unlevered beta using Hamada equation

Unlevered beta = Equity beta/(1+(1-t)*D/E) = 2.10/(1+(1-0.3)*1.4) = 1.0606

Equity beta for AMR would be Unlevered beta *(1+(1-t)*D/E)) = 1.0606*(1+(1-0.3)*0.42) = 1.37

Cost of equity for AMR = Rf + beta*(Rm-Rf) = 5.2% +1.37*(11.2%-5.2%) = 13.42%

AMR's equity cost of Capital = 13.42%

b. Weighted average cost of capital for AMR = 0.42/1.42 * 5.2%*(1-0.3) + 1/1.42 *13.42% = 10.53%

AMR's net value  = Cash flow next year/(r-g) = 15.4 Million/(0.1053-0.042) = 243.286 Million

AMR's share price = 243.286 Million/8 million = $30.41/share

AMR's share price = $30.41